OpenStudy (anonymous):

A trampoline company sells covers for several different sizes of trampolines. One trampoline has a diameter of 16 feet, and its cover sells for $180. The other trampoline has a diameter of 8 feet. How much should the company charge for the smaller trampoline cover, in dollars, if they want it to have the same price per square foot as the larger one?

OpenStudy (anonymous):

logic would force me to say, if it's half the size it should cost half as much.

Directrix (directrix):

Any two circles are similar. Assuming that the covers are circles, they are also similar with a scale factor of two to one or (2/1). The square of the scale factor is the ratio of the areas of the covers. (2/1) ^ 2 = 180/x where x is the cost of the smaller cover 4/1 = 180 / x 4x = 180 x = 45 dollars --> answer